OpenStudy (anonymous):

What is the derivative of h(t)=8sin(piet)+13t^1.3

OpenStudy (shamim):

is it\[h(t)=8 \pi \cos \pi t+13t ^{1.3}\]

OpenStudy (shamim):

if yes then i can try to solve it

OpenStudy (anonymous):

is the first post that you poseted your answer to the problem or are you asking ifthat is the problem

OpenStudy (shamim):

\[\frac{ dh }{dt }=8 \pi \cos \pi t+13 \times 1.3t ^{0.3}\]

OpenStudy (anonymous):

ok

OpenStudy (shamim):

did u catch my solution

OpenStudy (anonymous):

is it the last post that you pasted

OpenStudy (anonymous):

h(t)=8sin(pie t)+13 t^1.3 d/dt {h(t)}=8piecos(piet)+13(1.3)t^(1.3-1) h'(t)=8pie cos(pie t)+16.9t^0.3

OpenStudy (shamim):

do u need any more explanation @surjit

OpenStudy (anonymous):

now how do I find h'(3)

OpenStudy (anonymous):